Nasreen bought 3 m 20 cm cloth for her shirt and 2 m 5 cm cloth for her trouser. Find the total length of cloth bought by her.

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Cloth of shirt = 3 m 20 cm
Cloth of trouser = 2 m 5 cm
Total length of cloth is
=3.20 + 2.05
= 5.25
∴Total length of cloth bought by Nasreen is 5.25 m

Thanks
Shraddha Singh
3 years ago
Dear Student,  Cloth of shirt = 3 m 20 cm3 m 20 cm = (3 + 20/100)cm = 3.20 cm  Cloth of trouser = 2 m 5 cm2 m 5 cm = (2+5/100)cm = 2.05 cm★ Total length of cloth is= 3.20 + 2.05                                             = 5.25 m